x6.3.5 E Question Help Find the area of the shaded region. The graph to the right depicts IQ scores of adults, and those scores are normally distributed with a mean of 100 and a standard deviation of 15 115 The area of the shaded region is(Round to four decimal places as needed.)
Answer #1

Solution:

Given that,

mean = \mu = 100

standard deviation = \sigma = 15

P(x > 115 )

= 1 - P (x < 115 )

= 1 - P ( x -  \mu/ \sigma ) < ( 115 - 100 / 15 )

= 1 - P ( z < 15 / 15 )

= 1 - P ( z < 1)

Using z table

= 1 - 0.8413

= 0.1587

The area of shaded region is  = 0.1587